3. A business owner decides to use a binomial distribution to solve one of his problems. He knows there is a 60 percent probability that he will sell one of his Red Oak floors to any customer who comes into his shop and views the wood. Out of the next 7 people who come into his shop, what is the probability that 3 people will buy a Red Oak floor? ___________ What is the probability that more than four people will buy the Red oak floor? ___________
Answer by stanbon(75887)

A business owner decides to use a binomial distribution to solve one of his problems. He knows there is a 60 percent probability that he will sell one of his Red Oak floors to any customer who comes into his shop and views the wood. Out of the next 7 people who come into his shop,
what is the probability that 3 people will buy a Red Oak floor? ___________
Using a TI calculator I get binompdf(7,0.6,3) = 0.1935..
-----------------------------------------------------------------
What is the probability that more than four people will buy the Red oak floor?
1 - binomcdf(7,0.6,4) = 0.4199...
----------------------------------------
Alternate solution:
P(4 < x < 7) = P(x=5) + P(x=6) + P(x=7)
= 7C5(0.6)^5(0.4)^2 + 7C6(0.6)^6(0.4) + 7C7(0.6)^7(0.4)^0
= 21*0.07776*0.16 + 7*0.04665*0.4 + 1*0.02799
= 0.26127 + 0.13062 + 0.02799
= 0.41988...
================
Cheers,
Stan H.
___________
Cheers,
Stan H.